Math 125I Winter 2003 Name:

Quiz Section:

Quiz #2

SHOW YOUR WORK

1. (5 points) Use the Midpoint Rule and n = 4 to approximate ∫_{1}^{6} tan \sqrt{x} dx. Round your answer to four decimal places.

2. (5 points) The speed of a jogger is given below at quarter-second intervals. Give upper and lower estimates of the distance he travels in the first one and a half seconds.

t (sec)  0 0.25 0.5 0.75 1 1.25 1.5
v (ft/sec) 0 5.7 9.3 11.4 10.8 10.6 11.2

Math 125I Winter 2003 Name:

Quiz Section:

Quiz #2

SHOW YOUR WORK

1. (5 points) Use the Midpoint Rule and n = 4 to approximate ∫_{1}^{6} ln(tan x) dx. Round your answer to four decimal places.

2. (5 points) The speed of a jogger is given below at quarter-second intervals. Give upper and lower estimates of the distance he travels in the first one and a half seconds.

t (sec)  0 0.25 0.5 0.75 1 1.25 1.5
v (ft/sec) 0 5.6 9.2 11.5 10.7 10.8 11.3